a) What is the mass percent of hydrogen in sulfuric acid?

b) What is the mas of iron in iron(III) Chlorate?
Answer #1

1)

H2SO4

mass of H = 2*1 = 2

total mass = 98

% of H = 2/98*100 = 2.0408 % of H

b)

FeCl3

MW of Fe = 55.5

MW of FeCl3 = 162.2

% Fe = mass o fFe/ total mass * 100 = 55.5/162.2*100 = 34.2170 %